# Limits to scalable evaluation at the frontier: LLM as Judge won’t beat twice the data

### v2

@@ C.6 Proposition 2
$bp=b+b-bp,$
i.e. $p=1$. With this, the constraint of $\operatorname*{\mathrm{AG}}=r$ turns into $b+q(1-b)=r$. The left side is clearly continuous and monotonous in $q$ and inserting the extreme values of zero and one, we obtain $r=b$ and $r=1$, such that we can find a solution for all intermediate values of $r$.
@@
$2q-2bq+2b=2,$
which is achieved at $q=1$. In that case, the constraint of $\operatorname*{\mathrm{AG}}=r$ turns into $bp+(1-b)=r$ with extreme values again lying at $1-b=r$ and $1=r$. ∎
@@ C.8 Proposition 8
We drop the parameters’ dependence on $m$ and set $q=1-b$ and $p=b$. Then $\operatorname*{\mathrm{JB}}=b(1-b)-(1-b)b=0.$ At the same time, as there is no judge bias, $\operatorname*{\mathbb{E}}\tilde{s}(m)=b=p$, such that $\rho(s(m),\tilde{s}(m))^{2}=0$. Lastly $\operatorname*{\mathrm{AG}}=b^{2}+(1-b)^{2}$, which can attain any value between $0.5$ and $1$. ∎

### v3

@@ C.6 Proposition 2
bp=b+bbp,bp=b+b-bp,
i.e. $p=1$. With this, the constraint of $\operatorname*{\mathrm{AG}}=r$ turns into b+q(1b)=rb+q(1-b)=r. The left side is clearly continuous and monotonous in $q$ and inserting the extreme values of zero and one, we obtain $r=b$ and $r=1$, such that we can find a solution for all intermediate values of $r$.
@@
2q2bq+2b=2,2q-2bq+2b=2,
which is achieved at $q=1$. In that case, the constraint of $\operatorname*{\mathrm{AG}}=r$ turns into bp+(1b)=rbp+(1-b)=r with extreme values again lying at 1b=r1-b=r and $1=r$. ∎
@@ C.8 Proposition 8
We drop the parameters’ dependence on $m$ and set q=1bq=1-b and $p=b$. Then $\operatorname*{\mathrm{JB}}=b(1-b)-(1-b)b=0.$ At the same time, as there is no judge bias, $\operatorname*{\mathbb{E}}\tilde{s}(m)=b=p$, such that $\rho(s(m),\tilde{s}(m))^{2}=0$. Lastly $\operatorname*{\mathrm{AG}}=b^{2}+(1-b)^{2}$, which can attain any value between $0.5$ and $1$. ∎